Use a standard algorithm to calculate: 5−2/3 -3/5
1
Expert's answer
2020-09-07T19:44:02-0400

By taking LCM of the given equation we get,


"5- \frac {2}{3}- \frac{3}{5} = 5* \frac{15}{15}- \frac {2*5}{3*5}- \frac{3*3}{5*3}"


"=\frac {75}{15}-\frac {10}{15}-\frac {9}{15}"

"= \frac{75-10-9}{15}= \frac{56}{15}"


Hence,

"5- \frac {2}{3}- \frac{3}{5} = \frac{56}{15} = 3.73333"